The freezing point of benzene, C6H6, is 5.500 °C at 1 atmosphere.  Kf(benzene) = 5.12 °C/m

In a laboratory experiment, students synthesized a new compound and found that when 12.69 grams of the compound were dissolved in 242.9 grams of benzene, the solution began to freeze at 4.745 °C. The compound was also found to be nonvolatile and a non-electrolyte.  

What is the molecular weight they determined for this compound?  

__?__ g/mol
Answer #1

Δ Tf = 5.500 - 4.745 = 0.755 oC

use:

Δ Tf = Kf*mb

0.755 = 5.12 *mb

mb= 0.1475 molal

m(solvent)= 242.9 g

= 0.2429 kg

use:

number of mol,

n = Molality * mass of solvent in Kg

= (0.1475 mol/Kg)*(0.2429 Kg)

= 3.582*10^-2 mol

mass(solute)= 12.69 g

use:

number of mol = mass / molar mass

3.582*10^-2 mol = (12.69 g)/molar mass

molar mass = 3.543*10^2 g/mol

Answer: 354 g/mol